OpenStudy (anonymous):

A farmer has cows and ducks truck. There are only these creatures on the truck. Altogether there are 15 heads and 44 legs on the truck. How many cows and ducks are on the truck?

OpenStudy (anonymous):

Plzzzzzz help

OpenStudy (anonymous):

@Alecia0806 @aliboo997 @aum @algebra_student @BritBratt13 @blahhedup101 @blahhedup101 @barbiegirl12 @Chuck_O_Beast @vipcarlos1

OpenStudy (anonymous):

x is the cows y is the ducks x+y = 15 cows have 4 legs ducks have 2 4x+2y=44 multiply the first equation by 2 2x+2y =30 4x+2y=44 -2x=-14 divide each side by -2 x=7 plug x=7 into the first equation, x+y=15 7+y=15 y=8 so there are 7 cows and 8 ducks

OpenStudy (anonymous):

sorry i just saw this ... i hope i helped and i wasnt to late lol